how much of an alloy that is 30 percent should be mixed with 600 ounces of an alloy that is 70 percent copper in order to get an alloy that is 50 percent copper
Answer by stanbon(75887)

how much of an alloy that is 30 percent should be mixed with 600 ounces of an alloy that is 70 percent copper in order to get an alloy that is 50 percent copper
-----
Equation:
active + active = active
0.30x + 0.70*600 = 0.50(x+600)
---------
30x + 70*600 = 50x + 50*600
20x = 20*600
x = 600 ounces (amt. of 30% alloy needed)
======================
Cheers,
Stan H.